# DiEmo-TTS: Disentangled Emotion Representations via Self-Supervised Distillation for Cross-Speaker Emotion Transfer in Text-to-Speech

### v1

@@ 1 Introduction
Reference-based style transfer is commonly used in cross-speaker emotion synthesis, but it requires disentangling speaker information from emotion embeddings to prevent residual speaker traits from affecting the target timbre. However, certain prosodic features are inherently associated with the speaker’s identity, making complete disentanglement challenging. Therefore, an effective disentangling method is crucial to adequately separate speaker identity from emotion-related prosodic features, ensuring clear and accurate emotional transfer without compromising the target speaker’s timbre. Researchers typically implement the disentanglement method via explicit speaker labels, such as the gradient reversal layer (GRL) [8], as employed in [2, 3, 9, 10]. However, using explicit labels for disentanglement introduces a trade-off between preserving emotional information and separating speaker identity, making hyperparameter optimization challenging and leading to synthesis quality [4, 9]. Vector quantization (VQ) [11] offers an alternative approach to separating information without relying on explicit labels. Although VQ is effective in separating information without explicit labels [3, 12], it often results in unintended information loss and requires complex optimization to balance compression with reconstruction quality. To address these issues, [4] proposed an orthogonal loss to compensate for the embedding of emotion for the loss of emotional information caused by the disentanglement of speaker information [13]. Although this approach disentangles speaker information, it relies on multiple complex loss functions, making the optimization process overly complex. Furthermore, due to relying on GRL for disentanglement, its effectiveness is inherently limited between feature separation and synthesis quality, limiting its ability to fully preserve both speaker identity and emotional expressiveness in the generated speech.
@@
To overcome these limitations, we propose DiEmo-TTS, a novel approach for cross-speaker emotion disentangling and transfer. Building on the success of self-supervised learning in speech synthesis [15, 16], our method employs teacher-student emotion encoders trained using a distillation with no labels (DINO) strategy [17]. To minimize emotional information loss during disentanglement and enhance emotional expressiveness in synthetic speech, we introduce cluster-driven sampling for DINO. Additionally, style-adaptive condition modules enhance adaptability to emotions and speaker characteristics. Extensive experiments demonstrate the effectiveness of our method in improving emotional expressiveness while preserving the target speaker’s identity. Our audio samples are publicly available [^1] https://choddeok.github.io/DiEmo-TTS-Demo/.
@@ 3.1 Experimental setup
We use the emotional speech dataset (ESD) [29], which consists of about 350 parallel utterances spoken by 10 English speakers with five emotional states. Following the prescribed data partitioning criteria, we trained ten speakers in total. However, for the cross-speaker emotion transfer scenario, only two target speakers for inference, a male (“0013”) and a female (“0019”), were trained exclusively on neutral utterances. Moreover, we used the MSP-Podcast corpus dataset [30], which contains emotional speech data labeled with emotion categories and emotional attribute values. We employ the AdamW optimizer with hyperparameters $\beta_{1}=0.9$ and $\beta_{2}=0.98$, and configure the learning rate for the TTS system training at $5\times 10^{-4}$. The training process of the TTS module was conducted over approximately 24 hours on a single NVIDIA RTX A6000 GPU. For the audio synthesis in our experiments, we utilize the official implementation of BigVGAN [31], along with its pre-trained model.
@@ 3.6 Ablation study
This paper introduces an emotion disentangling method as a key extension to the traditional DINO method. To evaluate its effectiveness, we conduct an ablation study. As shown in Table 3.3, our model significantly improves performance, explained as follows: 1) “w/o DCT” removes the dual conditioning transformer block, replacing it with concatenation conditioning. Unlike simple style transfer, integrating multiple speech features better balances emotion and speaker characteristics, enhancing speech quality. 2) “w/o CDS” and “w/o CS” denote the removal of cluster-driven sampling, where random cropping is applied to a single sampled utterance, and the exclusion of cosine similarity loss, respectively. The removal of CDS and CS causes the emotion encoder to focus solely on emotional information rather than achieving proper disentangling, resulting in poorer overall performance and higher emotion-related metrics. 3) “w/o IP” replaces formant-based information perturbation with noise augmentation using the MUSAN [37] and RIR [38] datasets. While background noise or reverberation adds distortion, formant perturbation preserves emotion while perturbing speaker identity, improving disentanglement in cross-speaker emotion transfer. These results confirm that the proposed method effectively disentangles emotions and enables the TTS model to transfer multiple emotional and speaker styles harmoniously.
@@ 4 Conclusion
We present DiEmo-TTS, a system for cross-speaker emotion transfer that employs a novel self-supervised distillation method to minimize emotional information loss while preserving speaker identity. By leveraging a self-supervised distillation method with cluster-based disentangling, our approach effectively minimizes emotional information loss without relying on explicit emotion labels. Additionally, our style-adaptive conditioning enhances expressiveness and stability, producing natural and balanced speech. Experimental results demonstrate that the proposed method represents a new state-of-the-art in learning speaker-irrelevant emotion embeddings, excelling in expressiveness, naturalness, and speaker identity preservation. Despite its advancements, the proposed method does contain the limitation of relying on emotional dimension pseudo-labels, which originate from a predictor trained on labeled data. Future work will explore unsupervised methods to reduce dependence on pre-trained predictors and enhance the model’s ability to handle unseen speakers, enabling more robust cross-speaker emotion transfer. These advancements will drive more robust and scalable cross-speaker emotion transfer, enabling expressive and natural speech synthesis across diverse datasets.

### v2

@@ 1 Introduction
Reference-based style transfer is commonly used in cross-speaker emotion synthesis, but it requires disentangling speaker information from emotion embeddings to prevent residual speaker traits from affecting the target timbre. However, certain prosodic features are inherently associated with the speaker's identity, making complete disentanglement challenging. Therefore, an effective disentangling method is crucial to adequately separate speaker identity from emotion-related prosodic features, ensuring clear and accurate emotional transfer without compromising the target speaker's timbre. Researchers typically implement the disentanglement method via explicit speaker labels, such as the gradient reversal layer (GRL) [8], as employed in [2, 3, 9, 10]. However, using explicit labels for disentanglement introduces a trade-off between preserving emotional information and separating speaker identity, making hyperparameter optimization challenging and leading to synthesis quality [4, 9]. Vector quantization (VQ) [11] offers an alternative approach to separating information without relying on explicit labels. Although VQ is effective in separating information without explicit labels [3, 12], it often results in unintended information loss and requires complex optimization to balance compression with reconstruction quality. To address these issues, [4] proposed an orthogonal loss to compensate for the embedding of emotion for the loss of emotional information caused by the disentanglement of speaker information [13]. Although this approach disentangles speaker information, it relies on multiple complex loss functions, making the optimization process overly complex. Furthermore, due to relying on GRL for disentanglement, its effectiveness is inherently limited between feature separation and synthesis quality, limiting its ability to fully preserve both speaker identity and emotional expressiveness in the generated speech.
@@
To overcome these limitations, we propose DiEmo-TTS, a novel approach for cross-speaker emotion disentangling and transfer. Building on the success of self-supervised learning in speech synthesis [15, 16], our method employs teacher-student emotion encoders trained using a distillation with no labels (DINO) strategy [17]. To minimize emotional information loss during disentanglement and enhance emotional expressiveness in synthetic speech, we introduce cluster-driven sampling for DINO. Additionally, style-adaptive condition modules enhance adaptability to emotions and speaker characteristics. Extensive experiments demonstrate the effectiveness of our method in improving emotional expressiveness while preserving the target speaker's identity. Our audio samples are publicly available [^1] https://choddeok.github.io/DiEmo-TTS-Demo/.
@@ 3.1 Experimental setup
We use the emotional speech dataset (ESD) [29], which consists of about 350 parallel utterances spoken by 10 English speakers with five emotional states. Following the prescribed data partitioning criteria, we trained ten speakers in total. However, for the cross-speaker emotion transfer scenario, only two target speakers for inference, a male (``0013'') and a female (``0019''), were trained exclusively on neutral utterances. Moreover, we used the MSP-Podcast corpus dataset [30], which contains emotional speech data labeled with emotion categories and emotional attribute values. We employ the AdamW optimizer with hyperparameters $\beta_{1}=0.9$ and $\beta_{2}=0.98$, and configure the learning rate for the TTS system training at $5\times 10^{-4}$. The training process of the TTS module was conducted over approximately 24 hours on a single NVIDIA RTX A6000 GPU. For the audio synthesis in our experiments, we utilize the official implementation of BigVGAN [31], along with its pre-trained model.
@@ 3.6 Ablation study
This paper introduces an emotion disentangling method as a key extension to the traditional DINO method. To evaluate its effectiveness, we conduct an ablation study. As shown in Table 2, our model significantly improves performance, explained as follows: 1) ``w/o DCT'' removes the dual conditioning transformer block, replacing it with concatenation conditioning. Unlike simple style transfer, integrating multiple speech features better balances emotion and speaker characteristics, enhancing speech quality. 2) ``w/o CDS'' and ``w/o CS'' denote the removal of cluster-driven sampling, where random cropping is applied to a single sampled utterance, and the exclusion of cosine similarity loss, respectively. The removal of CDS and CS causes the emotion encoder to focus solely on emotional information rather than achieving proper disentangling, resulting in poorer overall performance and higher emotion-related metrics. 3) ``w/o IP'' replaces formant-based information perturbation with noise augmentation using the MUSAN [37] and RIR [38] datasets. While background noise or reverberation adds distortion, formant perturbation preserves emotion while perturbing speaker identity, improving disentanglement in cross-speaker emotion transfer. These results confirm that the proposed method effectively disentangles emotions and enables the TTS model to transfer multiple emotional and speaker styles harmoniously.
@@ 4 Conclusion
We present DiEmo-TTS, a system for cross-speaker emotion transfer that employs a novel self-supervised distillation method to minimize emotional information loss while preserving speaker identity. By leveraging a self-supervised distillation method with cluster-based disentangling, our approach effectively minimizes emotional information loss without relying on explicit emotion labels. Additionally, our style-adaptive conditioning enhances expressiveness and stability, producing natural and balanced speech. Experimental results demonstrate that the proposed method represents a new state-of-the-art in learning speaker-irrelevant emotion embeddings, excelling in expressiveness, naturalness, and speaker identity preservation. Despite its advancements, the proposed method does contain the limitation of relying on emotional dimension pseudo-labels, which originate from a predictor trained on labeled data. Future work will explore unsupervised methods to reduce dependence on pre-trained predictors and enhance the model's ability to handle unseen speakers, enabling more robust cross-speaker emotion transfer. These advancements will drive more robust and scalable cross-speaker emotion transfer, enabling expressive and natural speech synthesis across diverse datasets.